OpenStudy (anonymous):

Kelly tells you that when variables are in the denominator, the equation becomes unsolvable. "There is a value for x that makes the denominator zero, and you can't divide by zero," Kelly explains. Using complete sentences, demonstrate to Kelly how the equation is still solvable.

OpenStudy (anonymous):

Without knowing what the equation is, it's not easy to say, so I'll use an example. Suppose you have \[\frac{1}{x}=1\] Obviously, \(x\) can't be 0, as that makes the left side undefined. Since \(x\not=0\), you can multiply both sides by \(x\): \[\frac{x}{x}=x\] then the left side reduces to get \[1=x\]

OpenStudy (anonymous):

imsooo confused

OpenStudy (anonymous):

Not sure how else to explain this, sorry...

OpenStudy (anonymous):

can you try it with this equation1/x + 1/(x)² = 2